2. order:cefaclor oral suspension 200mg po q8h for a child with otitis media weighs 33 pounds. the recommended dose is 40mg/kg/d in 3 divided dose. is the ordered dose safe?

8.order: animophylline 27mg/kg/day po divided in 6 doses. the strength of the label is 105mg/5ml. calculate how many millilitres of this bronchodilator you will administer to a child who weighs 45lbs.

need step by step process to find the answer
Answer #1

1.recommended dose is 40mg/kg/day

Weight 33pound14.96 ie, 15

Si, 40×14=560mg for 24 hours

For 3divided dose ie, 560/3=186.66ie, 190mg

2. Order is 27mg/kg/day

Child weight 45pounds, so 20.41kg ie, 20kg

So, 27×20=540mg in 6divided dose

So, 540/6=90mg

What we have is 105mg

What we want 90mg

Amount to be dissolved is 5ml

What we want Milliliter of aminophyllin= what we want/what we have×amount to be dissolved

90/105×5=4.28

4.3 ml to be administered.